I need help in solving the following equation. Thanks so much.
7[4x+2] + 6=4

Answer by checkley77(12844)

|4x+2| means the positive & the negative value of the term inside the bars.
7|4x+2|+6=4
28x+14+6=4
28x+20=4
28x=4-20
28x=-16
x=-16/28
x=-4/7 ans.
------------------
7*-(4x+2)+6=4
-28x-14+6=4
-28x=4+14-6
-28x=12
x=12/-28
x=-3/7 ans.